OpenStudy (anonymous):

Please help with Alg 2 - Stat the answer choices are A) 0.3650 B) 0.7761 C) 0.2177 D) 0.7823

OpenStudy (anonymous):

You should avoid posting personal information on the internet. That being said, you can use z scores. The z score is how many standard deviations a value is from the mean. In this case, it would be:\[\frac{64.2-59}{6.7}=\frac{5.2}{6.7}\approx 0.77\] Knowing how many standard deviations past the mean you are can tell how much area is to the left of 59 by using a z table.